No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

B1_SD,B2_SD and B3_SD alarms due to low optical power

Publication Date:  2012-07-25 Views:  39 Downloads:  0
Issue Description
Equipment:- OSN 3500
Version:-  5.21.16.13
This NE is connected with other NE with SL16 cards.,
Alarm Information
There were B1_SD,B2_SD and B3_SD alarms on 13-SL16 card of OSN 3500.
Handling Process
After checked the received optical power at SL16 card and found that its received power was very low due to high fiber loss. When received power came to specified range then these alarms have cleared.
Root Cause
Null
Suggestions
Null

END